# CanadaFireSat: Toward high-resolution wildfire forecasting with multiple modalities

### v1

@@ 6 Conclusion
In this paper, we introduced CanadaFireSat, a comprehensive benchmark dataset for high-resolution wildfire forecasting over Canada from 2016 to 2023. CanadaFireSat was constructed to support multiple settings for model training: [Uncaptioned image]satellite images only, [Uncaptioned image]environmental predictors only, and [Uncaptioned image]satellite and environmental data. We demonstrated experimentally the potential of multi-modal learning for high-resolution wildfire forecasting on CanadaFireSat across two architectures: ResNet and ViT. Moreover, our experiments showed the importance of negative sampling in the evaluation of wildfire forecasting models. CanadaFireSat aims to accelerate research towards high-resolution monitoring of at-risk regions of interest to support wildfire management teams who are tasked with monitoring and protecting vast areas, such as the boreal ecosystem covering much of Canada. Results from this work demonstrate the feasibility of constructing future datasets like CanadaFireSat for other fire-prone landscapes, like the Pan-Arctic, Pan-boreal, and grassland and forest ecosystems of the Tropics. We hope this dataset and benchmark will foster research in this direction. To facilitate that, all codes, models, and CanadaFireSat are made publicly available on GitHub and HuggingFace.
@@ Appendix A Model Architectures
In this section, we illustrate the architectures used in the settings [Uncaptioned image]satellite images only and [Uncaptioned image]satellite and environmental data for the CNN-based models: Figure 12 and 13. We then show those used in the Transformer-based models: Figure 14 and 15.
@@ B.1 CNN Architecture Parameters
The model using only environmental predictora leverages inputs of dimension $H_{\text{mid}}=W_{\text{mid}}=32$ for mid-resolution data: MODIS, and $H_{\text{low}}=W_{\text{low}}=32$ for low-resolution data: ERA5, CEMS. MODIS data at $1$ km: LST Day is interpolated to $500$m to align with the rest of the MODIS inputs. Similarly, for the CEMS data: Fire Weather Index, and Drought Code, at 0.25°, is interpolated to $11.1$ km to align with ERA5-Land. We leverage the same set of environmental predictors as in the multi-modal setting, split into the two resolution groups. The temporal dimension of those inputs is $T_{\text{env}}=8$, the number of mid-resolution predictors is $N_{\text{mid}}=6$ including the day of the year, and the number of low-resolution predictors is $N_{\text{low}}=10$ including the day of the year (one more dimension than $N_{\text{env}}$ as we include the day of the year twice). As mentioned in Section 3.1, for the mid-resolution group we leverage $N_{S}=5$ multi-scale feature maps of dimensions: 64, 256, 512, 1024, and 2048, for the mid-resolution data. The last feature map of channel dimension 2048 is one-dimensional and is passed as such to an LSTM network for the temporal encoding. For the other four, we use independent ConvLSTM models. Those temporal encoders output feature maps of dimensions: 64, 128, 256, 512, and 1024, which are passed to a U-Net-like decoder and interpolated to the target size. The low-resolution inputs are encoded via a smaller network: ResNet-18, which outputs feature maps of channel dimension $D_{\text{low}}=512$, encoded temporally with a LSTM with of one layer to a dimension $D^{\prime}_{\text{low}}=64$, matching the channel dimension of the last feature map of the U-Net decoder. Both ResNet encoders are pre-trained on ImageNet. As for the other settings, the training is done with a target resolution of size $H_{\text{fire}}=W_{\text{fire}}=24$, and at test time the target resolution is $H_{\text{fire}}=W_{\text{fire}}=26$.
@@ B.2 ViT Architecture Parameters
For the environmental-only architecture, the input and target spatial dimensions and processing are identical to the CNN use case. The temporal dimension differs as we use $T_{\text{env}}=5$ for data augmentation. Both mid-resolution and low-resolution ViT-S encoders are randomly initialized and therefore do not use the LORA method for fine-tuning. In both encoders for the position embedding, attention, and projection, we use a dropout $d_{\text{env}}=0.2$ and a stochastic depth rate of $d_{\text{depth}}=0.1$. For mid-resolution inputs, the patch size is 2, and for low-resolution inputs, the patch size is 8. The temporal encoding of the mid-resolution feature map is identical to the one used for the satellite image time series, and the temporal encoding of the low-resolution data is done through a one-layer LSTM with both input and output channel dimensions $D_{\text{low}}=384$.
@@ B.3 CNN Training Parameters
The CNN models are trained using the combined weighted cross-entropy and dice loss. The positive class (fire) weight is $0.87$ and the negative class (no fire) is $0.13$, found experimentally. Training is run over 20 epochs with a batch size of 24 samples on a NVIDIA GeForce RTX 3080 Ti GPU. The scheduler for the learning rate follows a 2-epoch warm-up from the starting learning rate of $1e^{-7}$ to the base learning rate of $5e^{-6}$. Then the learning rate follows a cosine annealing of one cycle to the minimum learning rate of $1e^{-7}$ over the rest of the epochs. The optimizer used is ADAMW with a weight decay of $0.01$. During training, the augmentation pipeline first randomly crops the satellite input images to the training resolution, then resizes the images with a scale $s\in[0.9,1]$. The images are randomly flipped horizontally and vertically, and Gaussian noise injection is done with variance $\sigma^{2}\in[0.01,0.1]$. Finally, we randomly sample the satellite image time series to extract $T=5$ images (or pad when necessary). At test time, we center-crop the images to the required resolution and select the last $T=5$ samples. For the multi-modal training, the non-spatial environmental data is not augmented, while for the environment-only architecture, we apply random horizontal and vertical flipping and Gaussian noise injection, similarly to the satellite image time series. The missing values in the environmental predictors, mainly caused by the NDVI and EVI as they are 16-day composites, are replaced during training with the value $0.0$.
@@ B.4 ViT Training Parameters
Most of the ViT training parameters are the same as for the CNN models, except for the batch, which, despite also being 24, is accumulated across two steps of 12 for the ViT models. Moreover, during the training of the environmental-only use case, as we select $T_{\text{env}}=5$ time steps, it is also necessary to randomly sample across the total 8 available samples. Finally, at test time across all modalities, we use the native temporal length for each sample, 8 for the environmental data and a variable length for the satellite image time series. The processing of the missing values for the environmental predictors is the same as for the CNN-based architecture.
@@ Appendix C Ablation Study of the Impact of Satellite Image Time Series
In Table 5, we analyze the performance of the multi-modal models in setting [Uncaptioned image]satellite and environmental data with respect to the usage of time series. We compare our full multi-modal model using satellite image time series against a version using only the most recent image available before the prediction. In practice, for the CNN-based model, this impacts the number of parameters in the U-Net decoder as $D_{i}>D^{\prime}_{i}$. Table 5 shows the impact of SITS on model performance in setting [Uncaptioned image]satellite and environmental data, where, regardless of the architectures, the model performs best when presented with SITS rather than a single Sentinel-2 tile. As a consequence, we can hypothesize that dynamic factors directly linked to wildfire can be learned by the model from the temporal dimension of Sentinel-2.
@@ Appendix D Test Hard Analysis
Figure 16 demonstrates the domain shift between the Val and the Test set, as the evolution of the F1 score with the probability threshold is centered around 0.5 for the Val set, presenting a normal behavior while being shifted towards a smaller threshold value for the Test set. As a consequence, the metrics in Table 4 might overestimate the model performance on the test set due to the extreme fire patterns during this year. For this purpose, we constructed the adversarial set named Test Hard for the year 2023 as described in Section 4. Figure 16 also shows the delta in performance between Test and Test Hard and the centering of the maximum value for Test Hard closer to the 0.5 threshold, representing a better alignment with the model behavior on the Val set.
@@
Figure 17 presents the change in land cover distribution for the negative samples between the two sets: Test and Test Hard, with respect to the positive samples. The stratification sampling, done in Test Hard, aligns better the categorical distributions for the negative and positive populations.
@@ Appendix E Deployment at Scale: Second Case Study
In Figure 18, we present another case study for our CNN-based model in setting [Uncaptioned image]satellite and environmental data. This example presents a large wildfire in British Columbia occurring on the $2023/07/01$, displayed over an RGB composite of a Sentinel-2 image of $16\>\text{km}\times 22\>\text{km}$. Comparing our predictions with the positive samples covering this region of interest, we observe how well the model delineates the urban interface on the left side of the wildfire and the rough approximation of its boundaries on the right side of the fire. However, we can also see at the top of the Sentinel-2 image that the model overestimates the wildfire extent, similarly to what we observed in 11.
@@
Figure 18: Row 1 Sentinel-2 tile from $2023/06/06$ of size $16\>\text{km}\times 22\>\text{km}$ before a large wildfire in British Columbia. Row 2 Fire polygons for the large wildfire on $2023/07/01$ over the same tile. Row 3 Binary model predictions (in red) over the $2.64\>\text{km}\times 2.64\>\text{km}$ center-cropped positive samples in black.

### v2

@@ 6 Conclusion
In this paper, we introduced CanadaFireSat, a comprehensive benchmark dataset for high-resolution wildfire forecasting over Canada from 2016 to 2023. CanadaFireSat was constructed to support multiple settings for model training: [Uncaptioned image]satellite images only, [Uncaptioned image]environmental predictors only, and [Uncaptioned image]satellite and environmental data. We demonstrated experimentally the potential of multi-modal learning for high-resolution wildfire forecasting on CanadaFireSat across two architectures: ResNet and ViT. Moreover, our experiments showed the importance of negative sampling in the evaluation of wildfire forecasting models. CanadaFireSat aims to accelerate research towards high-resolution monitoring of at-risk regions of interest to support wildfire management teams who are tasked with monitoring and protecting vast areas, such as the boreal ecosystem covering much of Canada. Results from this work demonstrate the feasibility of constructing future datasets like CanadaFireSat for other fire-prone landscapes where high-resolution fire polygons are available, like the Pan-Arctic, Pan-boreal, and grassland and forest ecosystems of the Tropics, since all input variables are globally available and open-access, even though certain fire regimes might require other high-resolution sensors, as seen for peatland fires. We hope this dataset will foster research in this direction. To facilitate that, all codes, models, and CanadaFireSat are made publicly available on GitHub and HuggingFace.
@@ Appendix A Model Architectures
In this section, we illustrate the architectures used in the settings [Uncaptioned image]satellite images only (Figure 12) and [Uncaptioned image]satellite and environmental data (Figure 13) for the CNN-based models. We then show those used in the Transformer-based models in Figures 14 and 15, respectively.
@@ B.1 CNN Architecture Parameters
The model using only environmental predictors leverages inputs of dimension $H_{\text{mid}}=W_{\text{mid}}=32$ for mid-resolution data (MODIS), and $H_{\text{low}}=W_{\text{low}}=32$ for low-resolution data (ERA5, CEMS). MODIS data at $1$ km: LST Day is interpolated to $500$ m to align with the rest of the MODIS inputs. Similarly, for the CEMS data, the Fire Weather Index and Drought Code, originally at 0.250.25°, are interpolated to $11.1$ km to align with ERA5-Land. We leverage the same set of environmental predictors as in the multi-modal setting, split into the two resolution groups. The temporal dimension of those inputs is $T_{\text{env}}=8$, the number of mid-resolution predictors is $N_{\text{mid}}=6$ including the day of the year, and the number of low-resolution predictors is $N_{\text{low}}=10$ including the day of the year (one more dimension than $N_{\text{env}}$, as we include the day of the year twice). As mentioned in Section 3.1, for the mid-resolution group we leverage $N_{S}=5$ multi-scale feature maps of dimensions: $64$, $256$, $512$, $1024$, and $2048$, for the mid-resolution data. The last feature map of channel dimension $2048$ is one-dimensional and is passed to an LSTM network for the temporal encoding. For the other four, we use independent ConvLSTM models. Those temporal encoders output feature maps of dimensions $64$, $128$, $256$, $512$, and $1024$, which are passed to a U-Net-like decoder and interpolated to the target size. The low-resolution inputs are encoded via a smaller network: ResNet-18, which outputs feature maps of channel dimension $D_{\text{low}}=512$, encoded temporally with a LSTM with of one layer to a dimension $D^{\prime}_{\text{low}}=64$, matching the channel dimension of the last feature map of the U-Net decoder. Both ResNet encoders are pre-trained on ImageNet. As for the other settings, the training is done with a target resolution of size $H_{\text{fire}}=W_{\text{fire}}=24$, and at test time the target resolution is $H_{\text{fire}}=W_{\text{fire}}=26$.
@@ B.2 ViT Architecture Parameters
For the environmental-only architecture, the input and target spatial dimensions and processing are identical to the CNN use case. The temporal dimension differs as we use $T_{\text{env}}=5$ for data augmentation. Both mid-resolution and low-resolution ViT-S encoders are randomly initialized and therefore do not use the LORA method for fine-tuning. In both encoders, for the position embedding, attention, and projection, we use a dropout rate $d_{\text{env}}=0.2$ and a stochastic depth rate of $d_{\text{depth}}=0.1$. For mid-resolution inputs, the patch size is $2$, and for low-resolution inputs, the patch size is $8$. The temporal encoding of the mid-resolution feature map is identical to the one used for the satellite image time series, and the temporal encoding of the low-resolution data is done through a one-layer LSTM with both input and output channel dimensions $D_{\text{low}}=384$.
@@ B.3 CNN Training Parameters
The CNN models are trained using the combined weighted cross-entropy and dice loss. The positive class (fire) weight is $0.87$ and the negative class (no fire) is $0.13$, found experimentally. Training is run over $20$ epochs with a batch size of $24$ samples on a NVIDIA GeForce RTX 3080 Ti GPU. The scheduler for the learning rate follows a $2$-epoch warm-up from the starting learning rate of $1e^{-7}$ to the base learning rate of $5e^{-6}$. Then the learning rate follows a cosine annealing of one cycle to the minimum learning rate of $1e^{-7}$ over the rest of the epochs. The optimizer used is ADAMW with a weight decay of $0.01$. During training, the augmentation pipeline first randomly crops the satellite input images to the training resolution, then resizes the images with a scale $s\in[0.9,1]$. The images are randomly flipped horizontally and vertically, and Gaussian noise with variance $\sigma^{2}\in[0.01,0.1]$ is injected. Finally, we randomly sample the satellite image time series to extract $T=5$ images (or pad when necessary). At test time, we center-crop the images to the required resolution and select the last $T=5$ samples. For the multi-modal training, the non-spatial environmental data is not augmented, while for the environment-only architecture, we apply random horizontal and vertical flipping and Gaussian noise injection, similarly to the satellite image time series. The missing values in the environmental predictors, mainly caused by the NDVI and EVI as they are 16-day composites, are replaced during training with the value $0.0$.
@@ B.4 ViT Training Parameters
Most of the ViT training parameters are the same as for the CNN models, except for the batch, which, despite also being $24$, is accumulated across two steps of $12$ for the ViT models. Moreover, during the training of the environmental-only use case, as we select $T_{\text{env}}=5$ time steps, it is also necessary to randomly sample across the $8$ available samples. Finally, at test time across all modalities, we use the native temporal length for each sample, $8$ for the environmental data, and a variable length for the satellite image time series. The processing of the missing values for the environmental predictors is the same as for the CNN-based architecture.
@@ Appendix C Ablation Study of the Impact of Satellite Image Time Series
In Table 5, we analyze the performance of the multi-modal models in setting [Uncaptioned image]satellite and environmental data with respect to the usage of time series. We compare our full multi-modal model using satellite image time series against a version using only the most recent image available before the prediction. In practice, for the CNN-based model, this impacts the number of parameters in the U-Net decoder as $D_{i}>D^{\prime}_{i}$. Regardless of the architectures, the model performs best when presented with SITS rather than a single Sentinel-2 tile. As a consequence, we can hypothesize that dynamic factors directly linked to wildfire can be learned by the model from the temporal dimension of Sentinel-2.
@@ Appendix D Test Hard Analysis
Figure 16 demonstrates the domain shift between the Val and the Test set, as the evolution of the F1 score with the probability threshold is centered around 0.5 for the Val set, presenting a normal behavior while being shifted towards a smaller threshold value for the Test set. As a consequence, the metrics in Table 4 might overestimate the model performance on the test set due to the extreme fire patterns during this year. For this purpose, we constructed the adversarial set named Test Hard for the year 2023 as described in Section 2.1.2. Figure 16 also shows the delta in performance between Test and Test Hard: the centering of the maximum value for Test Hard is closer to the 0.5 threshold, representing a better alignment with the model behavior on the Val set.
@@
Figure 17 presents the change in land cover distribution for the negative samples between the two sets, Test and Test Hard, with respect to the positive samples. The stratification sampling done in Test Hard better aligns the categorical distributions for the negative and positive populations.
@@ Appendix E Deployment at Scale: Second Case Study
In Figure 18, we present another case study for our CNN-based model in setting [Uncaptioned image]satellite and environmental data. This example presents a large wildfire in Québec occurring on the $2023/07/05$, displayed over an RGB composite of a Sentinel-2 image of $14\;\text{km}\times 26\;\text{km}$. The predictions follow the same pattern as the actual wildfire, despite slightly overestimating its extent, as it can be seen on both sides of the Sentinel-2 tile, similarly to what we observed in Figure 11.
@@
Figure 18: Row 1 Sentinel-2 tile from $2023/06/28$ of size $14\>\text{km}\times 26\>\text{km}$ before a large wildfire in Québec. Row 2 Fire polygons for the large wildfire on $2023/07/05$ over the same tile. Row 3 Binary model predictions (in red) over the $2.64\>\text{km}\times 2.64\>\text{km}$ center-cropped positive samples (patches boundaries are outlined in black).